OpenStudy (anonymous):

A pair of equations is shown below: y = 7x − 8 y = 5x − 2 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer.

OpenStudy (onepieceftw):

For substitution solve on of the equations for a variable, say x For the first equation solving for x would look like this: \[\frac{ y+8 }{ 7 } = x\] The take that value for x (y+8/7) and substitute it for x in the second equation. \[y = 5(\frac{ y+8 }{ 7 }) - 2\] and now you can solve it.